A hydraulic system has two pistons of different diameter and uses a liquid of density p=850kg/m3. the smaller piston has a diameter of 4,9 cm and the mass of 1600 g and larger piston has a diameter of 14 cm and a mass of 2600 g. determine h, the height difference between the two pistons.


1
Expert's answer
2021-07-09T11:51:45-0400

"m_1g/A_1+\rho g h_1=m_2g/A_2+\rho g h_2\to"


"\Delta h=\frac{1}{\rho\cdot\pi}(\frac{4m_1}{d_1^2}-\frac{4m_2}{d_2^2})=\frac{1}{850\cdot3.14}(\frac{4\cdot 1.6}{0.049^2}-\frac{4\cdot 2.6}{0.14^2})=0.8\ (m)" . Answer